OpenStudy (anonymous):

what is the ratio of 8:5= ?% and how please break it down to the easiest terms please

OpenStudy (curry):

140%

OpenStudy (curry):

8 over 5 is 1.4

OpenStudy (curry):

so 1.4 = 140%

OpenStudy (dumbcow):

8:5 is same as 8/5 8/5 = 1.6 1.6 as percent, multiply by 100 160%

OpenStudy (anonymous):

the answer is 125% but how

OpenStudy (dumbcow):

i dunno how they got 125%

OpenStudy (anonymous):

i got the answers right here i just need to know the solution

OpenStudy (anonymous):

the answers are multiple choice and they are 75%, 150% 175% and 125% and the correct one is 125%

OpenStudy (dumbcow):

i showed you the solution, however it disagrees with what they want are you sure you wrote the question correctly?

OpenStudy (anonymous):

yes the ratio of 8:5=(?)%

OpenStudy (dumbcow):

well i can guarantee that 8/5 = 160%

OpenStudy (anonymous):

ok i trust u they could b wrong

OpenStudy (dumbcow):

yeah maybe or they are looking for something else, does this have to do with odds?

OpenStudy (anonymous):

3/5=(?)% *5/8

OpenStudy (dumbcow):

is this a different problem? ok 3/5 = x/100*(5/8) multiply by 8/5 on both sides (8/5)(3/5) = x/100 x/100 = 24/25 x = 2400/25 = 96